OpenStudy (anonymous):

Solve: 4(x − 1) − (x + 6) = 2

OpenStudy (anonymous):

i can solve it

OpenStudy (anonymous):

can i have a best response?

OpenStudy (anonymous):

4(x-1)-(x+6)=2 4x-4-x+6=2 3x+6-4=2 3x+2=2 3x+2-2=2-2 (x=3) *you distribute the 4 (: